OpenStudy (anonymous):

If a gas is heated from 298.0 K to 398.0 K and the pressure is increased from 2.230 x 103 mm Hg to 4.560 x 103 mm Hg what final volume would result if the volume is allowed to change from an initial volume of 60.0 liters?

OpenStudy (kendricklamar2014):

So we can use the formula: P1V1/T1 = P2V2/T2 For the first problem, we can plug the numbers we know into the formula, so it becomes 1200x85/363 = 850x350/T2 102000/363 = 297500/T2 Inverse and rearrange to get T2 on the left and above the divisor line T2/297500 = 363/102000 Move the 297500 to the right of the equal sign T2 = 363x297500/102000 T2 = 1058.75 Second problem uses the same principle 2.230x10^3 x 60 / 298 = 4.560x10^3 x V2 / 398 133.800x10^3 / 298 = 4.560x10^3 x V2 / 398 Move the values from the right to the left to leave V2 on its own. At the same time, rotate everything so V2 is on the left. V2 = (133.800x10^3 x 398) / (4.560x10^3 x 298) V2 = 39.1884

OpenStudy (anonymous):

thank u

OpenStudy (kendricklamar2014):

Your Welcome